# Generative Enzyme Design Guided by Functionally Important Sites and Small-Molecule Substrates

### v1

@@ Abstract
Enzymes are genetically encoded biocatalysts capable of accelerating chemical reactions. How can we automatically design functional enzymes? In this paper, we propose EnzyGen, an approach to learn a unified model to design enzymes across all functional families. Our key idea is to generate an enzyme’s amino acid sequence and their three-dimensional (3D) coordinates based on functionally important sites and substrates corresponding to a desired catalytic function. These sites are automatically mined from enzyme databases. EnzyGen consists of a novel interleaving network of attention and neighborhood equivariant layers, which captures both long-range correlation in an entire protein sequence and local influence from nearest amino acids in 3D space. To learn the generative model, we devise a joint training objective, including a sequence generation loss, a position prediction loss and an enzyme-substrate interaction loss. We further construct EnzyBench, a dataset with 3157 enzyme families, covering all available enzymes within the protein data bank (PDB). Experimental results show that our EnzyGen consistently achieves the best performance across all 323 testing families, surpassing the best baseline by 10.79% in terms of substrate binding affinity. These findings demonstrate EnzyGen’s superior capability in designing well-folded and effective enzymes binding to specific substrates with high affinities.
@@ 1 Introduction
  • We conduct experiments to evaluate EnzyGen and prior functional protein design methods. These results show that enzymes designed by EnzyGen lead to the best average enzyme-substrate ESP score (Kroll et al., 2023) of 0.65, substrate binding affinity of -9.44, and AlphaFold2 pLDDT (Jumper et al., 2021) of 87.45 across 323 families. EnzyGen outperforms the best baseline by 10.79% in terms of substrate binding affinity.
@@ 2 Related Work
Figure 1: (a) EnzyGen architecture, consisting of an enzyme modeling module (left) and a substrate representation module (right). The enzyme modeling module aims to generate the enzyme sequence and backbone structure, and the substrate representation module targets at predicting if an enzyme can bind to a substrate. The dashed box in enzyme input denotes functionally important sites, while other sites need to be generated. [M] denotes mask token. “1.1.1.1" denotes the fourth-level enzyme class in EC Tree. (b) Neighborhood equivariant layer: neighborhood message update (in green), neighborhood coordinate update (in blue) and neighborhood node feature update (in red). Indexed selection is choosing $\boldsymbol{x}_{j}$ (or $\boldsymbol{h}_{j}$) where $j^{th}$ residue is in the $K$-nearest neighbors of $i^{th}$ residue.
@@ 3.5 Functionally Important Enzyme Site Discovery
Functional sites are often evolutionarily conserved in a given enzyme family (Tristem, 2000; Bickel et al., 2002). Following previous methods (Armon et al., 2001; Panchenko et al., 2004; Chakrabarti & Lanczycki, 2007; Bray et al., 2009), we leverage multiple sequence alignment (MSA) to mine evolutionarily conserved patterns, which are known as functionally or structurally important sites. Designing enzymes conditioning on these conserved patterns can facilitate the generation of functional enzymes. We employ the ClustalW2 method (Anderson et al., 2011) to perform MSAs for each enzyme family. We designate residues as functionally important sites when they surpass a prespecified $\tau$ identity threshold.
@@ 4 Experiments
Table 1: ESP score ($\uparrow$) evaluating the enzyme-substrate interaction function. IF denotes the inverse folding model ProteinMPNN. A ESP score of 0.6 or higher indicates a positive substrate binding. Avg denotes average.
@@
Table 2: Substrate binding affinity ($\downarrow$) (kcal/mol). IF denotes the inverse folding model ProteinMPNN. Avg denotes average.
@@
Table 3: pLDDT ($\uparrow$) predicted by AlphaFold2. IF denotes the inverse folding model ProteinMPNN. Avg denotes average.
@@ 4.1 EnzyBench Construction
We aim to evaluate the effectiveness of our EnzyGen across a diverse range of enzymes. To achieve this, we gather all enzymes from BRENDA (Schomburg et al., 2002), yielding a total of $101,974$ PDB entries. These entries are classified into $3,157$ fourth-level enzyme categories based on the EC Tree (Appendix Figure 4). It is worth noting that there is a total of $8,422$ fourth-level categories, and our focus is specifically on those with experimentally confirmed structures. We then conduct MSAs for each fourth-level category to identify functionally important sites, using an MSA threshold ($\tau$) of 30%. Given the relatively small size of fourth-level categories, we aggregate data belonging to the same third-level category in the EC Tree, resulting in $256$ third-level categories. Subsequently, we cluster the PDB entries with a sequence identity threshold of $50\%$ to prevent data leakage. For robust evaluation, if the size of a third-level category is below 400, all the data are included in the training set; otherwise, we randomly split $100$ entries with 50 for validation and 50 for testing, while the remaining entries are utilized for training. We also ensure that no PDB entries in the validation or test sets belong to the same cluster as those in the training set. We further collect $23,711$ experimentally confirmed $<$enzyme, substrate$>$ pairs from Kroll et al. (2023), ensuring enzymes without a corresponding substrate have a randomly sampled negative one during the training process. We also make sure that all entries in the test set have a corresponding substrate to ensure function evaluation. Consequently, for validation and test, we obtain 30 third-level categories across 323 fourth-level categories. Detailed data statistics are provided in Appendix C.
@@ 4.2 Experimental Setup
Implementation Details. We use 33 NAELs to model enzymes and 3 neighborhood equivariant layers for substrate representation module with node feature dimensionality set to 1280. As suggested by Ying et al. (2018), the number of graph layers typically ranges from 2 to 6. To prevent over-fitting and facilitate efficiency, we utilize 3 neighborhood equivariant sub-layers for enzyme modeling, comprising a modified NAEL with every 11 global attention sub-layers followed by one neighborhood equivariant sub-layer. The parameters of the global attention sub-layer are initialized with the released 650M ESM-2 (Lin et al., 2022b) parameters. The total parameter count is 714M. The hyperparameters $\lambda/2$ and $K$ are set to $1.0$ and $30$, respectively. The model undergoes training for $1,000,000$ steps using 8 NVIDIA RTX A6000600060006000 GPU cards. The batch size and learning rate are set to $8192$ tokens and $3$e-$4$ respectively. Sequences are decoded using the greedy decoding strategy. We incorporate four-level enzyme tags in the EC tree. For example, an enzyme in the category 1.1.1.1 should have four tags: 1, 1.1, 1.1.1, and $1.$
@@
Function Evaluation. We evaluate the enzyme-substrate interaction function by the following metrics: (1) Following previous work (Li et al., 2013; Vidal-Limon et al., 2022), we use binding affinity, calculated by Gnina (McNutt et al., 2021) to quantify the strength of interactions between the designed enzymes and their substrates. A lower binding affinity indicates stronger binding. (2) We also use ESP score, developed by Kroll et al. (2023) to assess the ability of the designed enzymes to bind their corresponding substrates. In their paper, ESP model predicts enzyme-substrate interaction with 91% accuracy across multiple benchmarks. The scale of ESP score is 0-1, and a higher ESP score indicates a stronger enzyme-substrate interaction. Additionally, to evaluate if the designed enzymes are well-folded, we compute (3) pLDDT using AlphaFold2.
@@ 4.3 Main Results
The ESP scores, substrate binding affinities, and pLDDT results are shown in Tables 1, 2 and 3. For clarity, we average the performance of fourth-level categories under the same third-level one, with detailed results for 323 families in Appendix D.1. The results for top-1, top-5 and top-10 candidate are also provided in Appendix D.2.
EnzyGen excels in designing enzymes that bind their respective substrates with high affinities. Table 1 and 2 highlight the ability of EnzyGen in achieving the highest average ESP score and substrate binding affinity across 323 families. Notably, EnzyGen obtains an average ESP score of $0.65$, surpassing the suggested enzyme-substrate interaction threshold of 0.6 by the ESP evaluator developer. These demonstrate that enzymes designed by our EnzyGen are able to bind their corresponding substrates with high affinities.
EnzyGen is able to design well-folded enzymes. According to the findings of Guo et al. (2022) and Binder et al. (2022), high pLDDT scores (e.g., > 80) indicate high confidence of the residue structure. EnzyGen attains an average pLDDT score of $87.45$ across 323 families. This demonstrates the capability of our model to design enzymes with stable folding, validating the value of incorporating structural information into the enzyme design process.
@@ 5 Analysis: Diving Deep into EnzyGen
Figure 3: (a) ESP scores of designed enzymes from new fourth-level classes or with new substrates. Dash line denotes median. (b) Fourth-level category embedding clustering. Case Study: (c) Complex of designed 1KAG (2.7.1.71) and substrate ATP(-4), with pLDDT=90.39, Uniprot blastp recovery rate = 58.5%, (d) Complex of designed 5L2P (3.1.1.2) and substrate paraoxon, with pLDDT=89.44, Uniprot blastp recovery rate = 49.4%. Both cases show polar contacts (hydrogen bonds) depicted in purple.
@@
Table 4: ESP score ($\uparrow$) of EnzyGen and EnzyGen-finetune. Avg denotes average.
@@ 5.6 Designing Novel Enzymes with EnzyGen
EnzyGen is able to design novel enzymes that bind to specific substrates. Figure 3 (c) and (d) showcase two designed enzymes, both achieving pLDDT approaching or exceeding 90 and substrate binding affinities below -10. Additionally, the Gnina docked complexes show polar contacts (hydrogen bonds) between the enzymes and substrates. This affirms the enzyme-substrate interaction function. We carry out blastp search of designed enzymes in Uniprot, yielding an amino acid identity of 58.5% and 49.4%, respectively, to the most similar enzyme. Together, these results show that our model is capable of designing novel enzymes with high substrate binding affinities. Additional cases are detailed in Appendix D.3, each exhibiting a pLDDT higher than 80, a binding affinity below -10, and a Uniprot blastp identity rate lower than 65%. Enzyme sequences for all analyzed cases are also available in Appendix Table 19.
@@ 6 Conclusion
This paper introduces EnzyGen, a unified generative model for designing functional enzymes across diverse families. EnzyGen simultaneously generates enzyme sequence and backbone structure guided by automatically identified functionally important sites and a given substrate. We employ a set of three losses to train EnzyGen. To comprehensively evaluate EnzyGen, we construct EnzyBench, a benchmark for enzyme design, including all available enzymes within PDB across 3157 enzyme families. Experimental results demonstrate EnzyGen’s capability to design well-folded enzymes with strong enzyme-substrate interaction functions. A potential limitation is the lack of wet-lab testing for designed enzymes, which is currently underway in our laboratory.
@@ 4.6.1 Phosphorus-oxygen lyases (only sub-subclass identified to date)
Dataset Third-Level Category Fourth-Level Category Size Average Length
Training 256 3157 98974 327
Validation 30 428 1500 331
Test 30 323 1500 328
Table 6: Data statistics for EnzyBench.

### v3

@@ Abstract
Enzymes are genetically encoded biocatalysts capable of accelerating chemical reactions. How can we automatically design functional enzymes? In this paper, we propose EnzyGen, an approach to learn a unified model to design enzymes across all functional families. Our key idea is to generate an enzyme’s amino acid sequence and their three-dimensional (3D) coordinates based on functionally important sites and substrates corresponding to a desired catalytic function. These sites are automatically mined from enzyme databases. EnzyGen consists of a novel interleaving network of attention and neighborhood equivariant layers, which captures both long-range correlation in an entire protein sequence and local influence from nearest amino acids in 3D space. To learn the generative model, we devise a joint training objective, including a sequence generation loss, a position prediction loss and an enzyme-substrate interaction loss. We further construct EnzyBench, a dataset with 3157 enzyme families, covering all available enzymes within the protein data bank (PDB). Experimental results show that our EnzyGen consistently achieves the best performance across all 323 testing families, surpassing the best baseline by 10.79% in terms of substrate binding affinity. These findings demonstrate EnzyGen’s superior capability in designing well-folded and effective enzymes binding to specific substrates with high affinities. The code, model and dataset are released at https://github.com/LeiLiLab/EnzyGen.
@@ 1 Introduction
  • We conduct experiments to evaluate EnzyGen and prior functional protein design methods. These results show that enzymes designed by EnzyGen lead to the best average enzyme-substrate ESP score (Kroll et al., 2023) of 0.65, substrate binding affinity of -9.44, and AlphaFold2 pLDDT (Jumper et al., 2021) of 87.45 across 323 families (BRENDA enzyme classification fourth-level categories). EnzyGen outperforms the best baseline by 10.79% in terms of substrate binding affinity.
@@ 2 Related Work
Figure 1: (a) EnzyGen architecture, consisting of an enzyme modeling module (left) and a substrate representation module (right). The enzyme modeling module aims to generate the enzyme sequence and backbone structure, and the substrate representation module targets at predicting if an enzyme can bind to a substrate. The dashed box in enzyme input denotes functionally important sites, while other sites need to be generated. [M] denotes mask token. “1.1.1.1" denotes the fourth-level enzyme class in the BRENDA enzyme classification (EC) tree. (b) Neighborhood equivariant layer: neighborhood message update (in green), neighborhood coordinate update (in blue) and neighborhood node feature update (in red). Indexed selection is choosing $\boldsymbol{x}_{j}$ (or $\boldsymbol{h}_{j}$) where $j^{th}$ residue is in the $K$-nearest neighbors of $i^{th}$ residue.
@@ 3.3 NAEL Neighborhood Equivariant Sub-Layer
Figure 2: Discoverying functionally important sites. Each row is a protein sequence in a same enzyme family (BRENDA fourth-level enzyme classification tree category). We use ClustalW2 to perform multiple sequence alignment and select common residuals above the identity threshold $\tau$. According to the aligned sequences, E, G and M are common in all the sequences therefore these are selected as important sites. In experiment, $\tau=30\%$.
@@ 3.5 Functionally Important Enzyme Site Discovery
Functional sites are often evolutionarily conserved in a given enzyme family (Tristem, 2000; Bickel et al., 2002). Following previous methods (Armon et al., 2001; Panchenko et al., 2004; Chakrabarti & Lanczycki, 2007; Bray et al., 2009), we leverage multiple sequence alignment (MSA) to mine evolutionarily conserved patterns, which are known as functionally or structurally important sites. Designing enzymes conditioning on these conserved patterns can facilitate the generation of functional enzymes. We employ the ClustalW2 method (Anderson et al., 2011) to perform multiple sequence alignment for all proteins in each enzyme family corresponding to BRENDA fourth-level EC tree category. We select the residues that occur at the same site in more than $\tau$ sequences. These residues are designated as functionally important sites. In practice, we set $\tau$ to be $30\%$. Figure 2 illustrate the procedure to select functionally important sites.
Dataset Training Validation Test
Third-Level Category 256 30 30
Fourth-Level Category 3157 428 323
Size 98974 1500 1500
Average Seq. Length 327 331 328
Table 1: EnzyBench statistics. The first two rows are third and fourth-level categories in BRENDA enzyme classification tree. Size indicates the number of protein entries in PDB under the categories.
@@ 4 Experiments
Table 2: ESP scores ($\uparrow$) of generated enzymes for 30 testing BRENDA EC categories. ESP evaluates the enzyme-substrate interaction function. Each number is the average ESP score of 50 testing cases for each category. IF denotes the inverse folding model ProteinMPNN. A ESP score of 0.6 or higher indicates a positive substrate binding. The last column (Avg) is the average across 30 categories. Notice that our EnzyGen outperforms all previous methods in ESP by a big margin.
@@
Table 3: Enzyme-substrate binding affinity ($\downarrow$) (kcal/mol) for 30 testing BRENDA EC categories, evaluated by a docking software Gnina. IF denotes the inverse folding model ProteinMPNN. The last column (Avg) is the average across 30 categories. Notice that our EnzyGen outperforms all previous methods in binding affinity by a big margin.
@@
Table 4: pLDDT ($\uparrow$) predicted by AlphaFold2. IF denotes the inverse folding model ProteinMPNN. The last column (Avg) is the average across 30 categories. Notice that our EnzyGen outperforms all previous methods in pLDDT. It shows EnzyGen’s enzyme can fold more stably.
@@ 4.1 EnzyBench Construction
We aim to evaluate the effectiveness of our EnzyGen across a diverse range of enzymes. To achieve this, we gather all enzymes from BRENDA (Schomburg et al., 2002), yielding a total of $101,974$ PDB entries. These entries are classified into $3,157$ fourth-level enzyme categories based on the EC Tree (Appendix Figure 5). It is noteworthy that there exist a total of $8,422$ fourth-level categories; however, our focus lies specifically on those with experimentally confirmed structures. This focus results in our dataset comprising $3,157$ fourth-level categories. We then conduct MSAs for each fourth-level category to identify functionally important sites, using an MSA threshold ($\tau$) of 30%. Given the relatively small size of fourth-level categories, we aggregate data belonging to the same third-level category in the EC Tree, resulting in $256$ third-level categories. Then we select 30 third-level categories for validation and testing, respectively including 428 and 323 fourth-level categories. For each of the 30 third-level categories, we we randomly split $100$ PDB entries with 50 for validation and 50 for testing, while the remaining entries are utilized for training. To prevent data leakage, we cluster the PDB entries with a sequence identity threshold of $50\%$. We ensure that no PDB entries in the validation or test sets belong to the same cluster as those in the training set. We further collect $23,711$ experimentally confirmed $<$enzyme, substrate$>$ pairs from Kroll et al. (2023), ensuring enzymes without a corresponding substrate have a randomly sampled negative one during the training process. We also make sure that all entries in the test set have a corresponding substrate to ensure function evaluation. Detailed data statistics are provided in Table 1 and Appendix C.
@@ 4.2 Experimental Setup
Implementation Details. We use 33 NAELs to model enzymes and 3 neighborhood equivariant layers for substrate representation module with node feature dimensionality set to 1280. As suggested by Ying et al. (2018), the number of graph layers typically ranges from 2 to 6. To prevent over-fitting and facilitate efficiency, we utilize 3 neighborhood equivariant sub-layers for enzyme modeling, comprising a modified NAEL with every 11 global attention sub-layers followed by one neighborhood equivariant sub-layer. The parameters of the global attention sub-layer are initialized with the released 650M ESM-2 (Lin et al., 2022b) parameters. The total parameter count is 714M. The hyperparameters $\lambda/2$ and $K$ are set to $1.0$ and $30$, respectively. The model undergoes training for $1,000,000$ steps using 8 NVIDIA RTX A6000600060006000 GPU cards, among which only the sequence generation and position prediction losses are adopted for the first $200,000$ steps. The batch size and learning rate are set to $8192$ tokens and $3$e-$4$ respectively. Sequences are decoded using the greedy decoding strategy. We incorporate four-level enzyme tags in the EC tree. For example, an enzyme in the category 1.1.1.1 should have four tags: 1, 1.1, 1.1.1, and $1.$
@@
Function Evaluation. We evaluate the enzyme-substrate interaction function by the following metrics: (1) Following previous work (Li et al., 2013; Vidal-Limon et al., 2022), we use enzyme-substrate binding affinity, calculated by Gnina (McNutt et al., 2021) to quantify the strength of interactions between the designed enzymes and their substrates. A lower binding affinity indicates stronger binding. (2) We also use ESP score, developed by Kroll et al. (2023) to assess the ability of the designed enzymes to bind their corresponding substrates. In their paper, ESP model predicts enzyme-substrate interaction with 91% accuracy across multiple benchmarks. The scale of ESP score is 0-1, and a higher ESP score indicates a stronger enzyme-substrate interaction. Additionally, to evaluate if the designed enzymes are well-folded, we compute (3) pLDDT using AlphaFold2.
@@ 4.3 Main Results
The ESP scores, substrate binding affinities, and pLDDT results are shown in Tables 2, 3 and 4. For clarity, we average the performance of fourth-level categories under the same third-level one, with detailed results for 323 families in Appendix D.1. The results for top-1, top-5 and top-10 candidate are also provided in Appendix D.2.
EnzyGen excels in designing enzymes that bind their respective substrates with high affinities. Table 2 and 3 highlight the ability of EnzyGen in achieving the highest average ESP score and substrate binding affinity across 323 families. Notably, EnzyGen obtains an average ESP score of $0.65$, surpassing the suggested enzyme-substrate interaction threshold of 0.6 from the ESP evaluator developer. In comparison, PROTSEED achieves 0.59, RFDiffusion+ProteinMPNN achieves 0.53, and ESM2+EGNN achieves 0.61 of average ESP score. EnzyGen’s ESP score is significantly higher than all previous baselines. This is further verified by the binding affinity scores. The enzymes generated by our EnzyGen exhibit stronger binding affinity to their corresponding substrates than all other methods across all 30 third-level testing EC categories. EnzyGen gains an average improvement of $-0.87$ in binding affinity over the previous best method (RFDiffusion+ProteinMPNN), which is significant. These results demonstrate that enzymes designed by our EnzyGen are able to bind their corresponding substrates with high affinities.
EnzyGen is able to design well-folded enzymes. According to the findings of Guo et al. (2022) and Binder et al. (2022), high pLDDT scores (e.g., > 80) indicate high confidence of the residue structure. EnzyGen attains an average pLDDT score of $87.45$ across 323 families. It outperforms all three baselines (PROTSEED, RFDiffusion+ProteinMPNN, ESM2+EGNN). This demonstrates the capability of our model to design enzymes with stable folding, validating the value of incorporating structural information into the enzyme design process.
@@ 5 Analysis: Diving Deep into EnzyGen
Figure 4: (a) ESP scores of designed enzymes from new fourth-level classes or with new substrates. Dash line denotes median. (b) Fourth-level category embedding clustering. Case Study: (c) Complex of designed 1KAG (2.7.1.71, catalyzing the specific phosphorylation of the 3-hydroxyl group of shikimic acid) and substrate ATP(-4), with pLDDT=90.39, Uniprot blastp recovery rate = 58.5%, (d) Complex of designed 5L2P (3.1.1.2, hydrolyzing various p-nitrophenyl phosphates, aromatic esters and p-nitrophenyl fatty acids) and substrate paraoxon, with pLDDT=89.44, Uniprot blastp recovery rate = 49.4%. Both cases show polar contacts (hydrogen bonds) depicted in purple.
@@
Table 5: ESP score ($\uparrow$) of EnzyGen and EnzyGen-finetune. Avg denotes average. Finetuning leads to improvement in ESP scores.
@@ 5.6 Designing Novel Enzymes with EnzyGen
EnzyGen is able to design novel enzymes that bind to specific substrates. Figure 4 (c) and (d) showcase two designed enzymes based on PDB 1KAG (catalyzing the specific phosphorylation of the 3-hydroxyl group of shikimic acid) and 5L2P (hydrolyzing various p-nitrophenyl phosphates, aromatic esters and p-nitrophenyl fatty acids) respectively. Both of the two designed enzymes achieve pLDDT approaching or exceeding 90 and substrate binding affinities below -10. Additionally, the Gnina docked complexes show polar contacts (hydrogen bonds) between the enzymes and substrates. This affirms the enzyme-substrate interaction function. We carry out blastp search of designed enzymes in Uniprot, yielding an amino acid identity of 58.5% and 49.4%, respectively, to the most similar enzyme. Together, these results show that our model is capable of designing novel enzymes with high substrate binding affinities. Additional cases are detailed in Appendix D.3, each exhibiting a pLDDT higher than 80, a binding affinity below -10, and a Uniprot blastp identity rate lower than 65%. Enzyme sequences for all analyzed cases are also available in Appendix Table 19.
6 Discussion
Enzymes, as genetically encoded biocatalysts, play a crucial role in accelerating chemical reactions and find extensive applications across various fields. Designing enzymes that can specifically bind to target substrates has always been a complex challenge. Our EnzyGen demonstrates a remarkable ability to design well-folded and effective enzymes that bind to specific substrates across diverse enzyme categories. Despite its impressive performance, certain limitations remain, which we will address in this section.
First, we collected enzyme data from the PDB, covering $3,157$ fourth-level enzyme classes. Although there are $8,422$ fourth-level classes in total, our dataset includes only a subset of these categories. Future work could focus on expanding this dataset by incorporating data from larger databases, such as Swiss-Prot and the AlphaFold structure database, to encompass a broader range of enzyme classes.
Second, we address the substrate-binding constraint using an enzyme-substrate binding prediction loss. While this provides a basic constraint, it is relatively weak concerning the enzyme’s catalytic functions. Future research could focus on designing enzymes that bind to a specific 3D substrate structure, potentially leading to the formation of a more effective enzyme-substrate complex.
@@ 7 Conclusion
This paper introduces EnzyGen, a unified generative model for designing functional enzymes across diverse families. EnzyGen simultaneously generates enzyme sequence and backbone structure guided by automatically identified functionally important sites and a given substrate. We employ a set of three losses to train EnzyGen. To comprehensively evaluate EnzyGen, we construct EnzyBench, a benchmark for enzyme design, including all available enzymes within PDB across 3157 enzyme families. Experimental results demonstrate EnzyGen’s capability to design well-folded enzymes with strong enzyme-substrate interaction functions.
Acknowledgements
This research is supported by the National Institutes of Health (R35GM147387 to Y.Y.), a seed grant from the NSF Molecule Maker Lab Institute (grant #2019897), the UC Santa Barbara Faculty Research Grant (to L.L.), and the generous support by Pittsburgh Supercomputing Center Neocortex grant. The authors thank the anonymous reviewers and Siqi Ouyang, Yuwei Yang, Yufei Song, Jielin Qiu, and Yujia Gao for their valuable comments.